The period T(time in seconds for one complete cycle) of a simple pendulum is realated to length L (in feet)of the pendulum for the formula 8T^2 = PIE^2L. If a child is on a swing with a 10- foot chain, then how long does it take to complete on cycle of the swing?

Answer by stanbon(75887)

The period T(time in seconds for one complete cycle) of a simple pendulum is related to length L (in feet) of the pendulum for the formula 8T^2 = PIE^2L.
-------------------------
If a child is on a swing with a 10-foot chain, then how long does it take to complete on cycle of the swing?
--------------------------
8T^2 = (pi)^(2*10)
8T^2 = 62.83...
T = sqrt[62.83/8] = 2.8 seconds
======================
Cheers,
Stan H.